OpenStudy (ajtran):

PLEASE CHECK MY ANSWER <3 ~ the diameter of the base of a right cone is 12m . The total surface area of the cone is 306^cm2. What is the length of the slant height ? 1) 45 cm 2) 36 cm 3 13.5 cm < my answer 4) 27 cm

OpenStudy (zootedyoungin):

How did you get that?

OpenStudy (ajtran):

AHHH I meant to say my answer was 45cm ;-; sorry!

OpenStudy (zootedyoungin):

You were right the first time

OpenStudy (ajtran):

But when I plugged it into SA = πr² + πrs I got 45?

OpenStudy (ajtran):

Ahh now im lost >~< I thought it was 13.5 < but my friend told me it was wrong and now I got 45

OpenStudy (sshayer):

diameter is 12 meters.

OpenStudy (sshayer):

or 12 cms

OpenStudy (sshayer):

i doubt ,it can not be meters.

OpenStudy (sshayer):

i consider diameter=12 cms r=12/2=6 \[surface~area=\pi r^2+\pi r l\] \[\pi *6^2+\pi*6*l=306,6*\pi \left(\pi+l \right)=306,\pi \left( \pi+l \right)=\frac{ 306 }{ 6 }=51\] \[\pi+l=51\times \frac{ 7 }{ 22 }=16.23,l=16.23-\pi=16.23-\frac{ 22 }{ 7 }=13.08\]

OpenStudy (sshayer):

l=slant height=13.08 cms approx.13 cms